Alyssa and Gabriel play the following game. Alyssa rolls a die. If she rolls a​ 1, 2, or​ 3, Gabriel gives Alyssa

​$5.

If Alyssa rolls a 4 or​ 5, Gabriel gives Alyssa

​$4.

​However, if Alyssa rolls a​ 6, she gives Gabriel

​$24.

​a) Determine​ Alyssa's expectation.

​b) Determine​ Gabriel's expectation.
Answer #1

Probability of getting {1,2 or 3}=3/6
Probability of getting {4 or 5}=2/6

Probabilityof getting 6= 1/6

Now,

Que#a) Determine​ Alyssa's expectation.

Solutions:

Expectation (X)=$5*prob{1,2,3}+ $4*prob(4, 5) -$24* prob(x=6)

= 5*3/6+4*2/6-24*1/6=-$1/6

Ans= -$1/6.                                      (loss)

Que(b)

b) Determine​ Gabriel's expectation.

Solutions:

Gabriel Expectation

= - (Alyssa's expectation

=$1/6.       (,profit)

Ans=$1/6

Thanks!